zarkam21:

A farmer wishes to enclose a pasture that is bordered on one side by a river (so one of the four sides won't require fencing). She has decided to create a rectangular shape for the area, and will use barbed wire to create the enclosure. There are 600 feet of wire available for this project, and she will use all the wire. What is the maximum area that can be enclosed by the fence? (Hint: Use this information to create a quadratic function for the area enclosed by the fence, then find the maximum of the function.)

Vocaloid:

you can express the perimeter as 2W + L = 600 (since there are only 3 sides of fence used) area = W*L use the first equation to re-write area in terms of one variable (L or W) and then solve for that variable

Vocaloid:

so, let's just solve for L L = 600 - 2W then area becomes W*L = W*(600-2W) then expand using foil and find the maximum using the formula we used before (x, or in this case, W = -b/(2a)

zarkam21:

so w=-600/(2*2)

Vocaloid:

b and a are not 600 and 2 keep in mind you have to re-write the equation in standard form to find a, b, c so expand W*(600-2W) using foil

zarkam21:

wait so what are the two binomials that i will use to foil

Vocaloid:

should be the distributive property actually

Vocaloid:

A(B-C) = A*B - A*C W*(600-2W) = ?

zarkam21:

600w-600(2)

Vocaloid:

W*2W = ?

zarkam21:

w^2

zarkam21:

:/

Vocaloid:

don't forget the 2 W*2W = 2*W*W = 2W^2

Vocaloid:

so the entire function is -2W^2 + 600W making W = -b/(2a) = -600/(2*(-2)) = 150 W = 150 now, find L using the formula from before L = 600 - 2W then after that, find area = L*W

zarkam21:

l=300

Vocaloid:

good, now find area using W*L

zarkam21:

300*150=45000

Vocaloid:

good, that's your answer, 45,000 ft ^2